OpenStudy (anonymous):

The intensity of light from a light source varies inversely as the square of the distance from the light source. Write an equation for the variation. Use k as the constant, I for the intensity of light, and d for the distance.

OpenStudy (anonymous):

intensity say I = k/ (distance d)^2

OpenStudy (anonymous):

i inversely proportional d^2 i=k/d^2